• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name mercado

    Log entry time 11:48:36 on August24,2009

    Entry number 285693

    Followups:

    keyword=More Lumi commisioning. Attempt at Oversampling.

    We took some Lumi data this morning. The beam was 20uA on a Tantalum target.
    
    During run 11546, the HVs were tuned to get a 200K signal out of our 18-bit 
    ADCs.
    
    Run 11547 is a good run, which lasted 15 minutes and had 2 beam trips for the 
    purposes of pedestal studies and calibration. For this run, the BCM gains were 
    LO and the highest signal BCM that didn't saturate was bcm4, with a signal of 
    ~57K channels.
    
    Run 11549 is an oversampling run (IT=6600, OS=1). The gain setting of the 
    18-bit board was changed from (3,0) to (3,1), which has a gain factor of 2. The 
    BCMs were set to HI gain, and the highest signal BCM for this run was BCM2, 
    with a signal of ~33K channels.
    
    Run 11550 is an oversampling run (IT=3300, OS=3). The gain setting ofthe 18-bit 
    board was changed to (3,3), which has a gain factor of about 3.9. BCMs 1, 2 and 
    4 where set to HI gain, BCM3 was not working while connected to channel 23_2, 
    so I switched it to channel 23_3, where ot produced a signal of ~39K channels 
    at LO gain. BCMs 3 and 4 both have decent signals on this run.
    
    NOTE: During this last run, I realized that the Injector crate had not been 
    responding to GreenMonster. This means that the Injector data, which includes 
    the helicity data, is probably nonsense, since the IT value for that timing 
    board remained the same throughout the runs. Because of this, PAN will probably 
    not be able to calculate proper asymmetries for the lumis for the oversampling 
    runs. The last two runs should be repeated ASAP.